# Cheaper and more noise-resilient quantum state preparation using eigenvector continuation

### v1

@@ Abstract
Subspace methods are powerful, noise-resilient methods that can effectively prepare ground states on quantum computers. The challenge is to get a subspace with a small condition number that spans the states of interest using minimal quantum resources. In this work, we will use eigenvector continuation (EC) to build a subspace from the low-lying states of a set of Hamiltonians. The basis vectors are prepared using truncated versions of standard state preparation methods such as imaginary time evolution (ITE) and adiabatic state preparation (ASP). By using these truncated methods combined with eigenvector continuation, we can directly improve upon them, obtaining more accurate ground state energies at a reduced cost. We use several spin systems to demonstrate convergence even when methods like ITE and ASP fail, such as ASP in the presence of level crossings and ITE with vanishing energy gaps. We also showcase the noise resilience of this approach beyond the gains already made by having a shallower quantum circuits. Our findings suggest that eigenvector continuation can be used to improve existing state preparation methods in the near term.
@@ I Introduction
Finding the ground state of non-integrable systems is one of the primary targets of studies in quantum mechanics. This goal is hindered by the exponential growth in Hilbert space dimension, which limits the use of classical devices for this purpose; the largest-scale simulations nevertheless are capped at a relatively modest ($\sim 20$) number of lattice sites for spin models [1, 2]. Within the setting of quantum computing, the problem is recast as finding an efficient way to prepare a desired quantum state. This problem is generally QMA-hard [3, 4, 5], i.e. difficult even for quantum computers; nevertheless, a significant research effort is being directed towards making inroads into this important problem, as quantum devices are believed to be able to approach this problem and perform much better than their classical counterparts [6, 7]. This promise is not yet realized, and problems such as barren optimization plateaus [8, 9, 10] plague the current methods. While fault tolerant algorithms that have provable convergence rates do exist, these are not yet realizable on near-term hardware, which is noisy and limited in its capabilities. Because of this reason, researchers are focusing on developing NISQ-friendly algorithms that can still perform well with hardware imperfections [11, 12, 13].
Several quantum state preparation algorithms can be categorized into these broad categories, including variational quantum algorithms (e.g, VQE, QAOA) [14, 15, 16, 17, 18, 19, 20], time evolution methods (e.g, ASP, QITE) [21, 22, 23, 24, 25], and subspace methods (e.g, Krylov subspace) [26, 27, 28, 29, 30, 31]. While all of them have some advantages and disadvantages, none of them are ideal for implementation on current hardware. Variational quantum algorithms use classical optimization-based and learning-based approaches with parameterized quantum circuits but face optimization landscape issues along with trainability, accuracy, and efficiency limitations [8]. Time evolution methods, and in particular adiabatic state preparation (ASP), provide an efficient mapping to a quantum circuit; however, ASP often requires a long evolution [32, 33, 23], an easily prepared starting state, and cannot be used in the presence of symmetry-protected level crossings. Imaginary time evolution [34] (ITE) guarantees convergence, but implementing non-unitary evolution in a quantum circuit is complicated. Both evolution-based algorithms tend to involve deep quantum circuit implementations; and as the system size and the corresponding Hilbert space grow, these algorithms require a concomitantly longer circuit to converge [35, 36, 37, 38], thus increasing the quantum resources required.
@@
In this work, we will improve on this idea by realizing that an exact ground state is not needed for EC to work; rather, obtaining a number of states with a decent overlap with the low energy manifold suffices. Here we will use real- and imaginary-time ground-state preparation techniques, although other methods could be used similarly to get the low-energy states. In all cases, an otherwise insufficiently deep (or truncated) circuit can be used. Getting low-energy states is much easier and more cost-effective than obtaining exact ground states; we will explore these ideas for time evolution methods in the next section. A similar idea of using the approximate state preparation for getting a basis for subspace diagonalization was explored in a recent work [49], where they use approximate VQE for preparing a subspace to solve quantum chemistry instances, but remain focused on a single Hamiltonian.
The outline of the paper is as follows: In section II, we will outline the method of eigenvector continuation. In section III, we will apply it using time evolution state preparation methods to get a low-energy subspace for the XY spin model. In section IV, we will extend the application of the method to address the issue of vanishing energy gaps using the time-evolution methods with an example of a model with level crossings, XY model, and model with changing degeneracies in the ground state, Kagome XXZ model. Section V discusses the noise resilience of the method by performing noisy simulations. Lastly, we conclude with a discussion in section VI.
@@ II Eigenvector Continuation
In this work, we still use a set of basis that that spans the low-energy subspace, but we propose using state preparation methods that are truncated before they reach the ground state. This gives us a low-energy state that is not necessarily an eigenstate or the ground state, rather, this is a superposition of a few low-energy states of the system. As we will show, getting these states from the small number of training points gets us enough overlap with all the ground states to be able to obtain reasonable ground states from the generalized eigenvalue problem.
@@
For the training Hamiltonians, we find approximate ground states, using any state preparation method, to obtain a set of low-energy vectors that will form the basis vectors for a subspace $(\phi_{t}(\theta)=\{\ket{\phi_{1}},\ket{\phi_{2}},\dots\ket{\phi_{k_{p}}}\})$. In this paper we have focused on time evolution methods for state preparation—ASP and ITE—but other methods such as variational approaches [49] can be used similarly.
Now that we have a subspace that spans low energy states, we project the target Hamiltonians $(H(\theta_{i}))$ onto this subspace and solve the generalized eigenvalue problem
@@
The lowest eigenvalue thus obtained gives the ground state energy of the target Hamiltonian. The overlap matrix $(\mathbb{S})$ of the basis vectors remains the same for all the target points, but the projected Hamiltonian ($\mathbb{H}$) needs to be calculated separately for each target point.
The GEP is solved by inverting the overlap matrix and solving the equation,
$\mathbb{S}^{-1}\mathbb{H}(\theta)\ket{\psi}=E\ket{\psi}$ (2)
This requires the overlap matrix to be invertible, which means the determinant of the matrix should not vanish. Because of this reason, the number of training states needs to be limited to prevent over-spanning of the subspace. We diagnose this issue with the condition number $(C_{n})$ of the overlap matrix $(\mathbb{S})$, which is defined as the ratio of the largest eigenvalue to the smallest eigenvalue of the matrix. For $\mathbb{S}$ to be invertible, $C_{n}$ should be small.
@@ III Improving Quantum state preparation methods using EC
In this work, we will use time evolution based methods, i.e. adiabatic state preparation (ASP) and imaginary time evolution (ITE), for getting our low-energy subspace. Both methods improve with additional resources, which for both of these means both a larger total evolution time and a smaller time step; we will focus on the former. To illustrate the use of EC to improve a truncated state preparation, we truncate at $\leq 10\%$ of the time evolution steps required to get reasonable ground state energies.
@@ III.0.1 Imaginary Time Evolution
Figure 2: Implementing EC with truncated ASP and ITE for a 1D chain XY model 5 site. (a) and (d) plots the ground energy spectrum, (b) and (e) plots corresponding ground state fidelity change. (d) and (f) plot RMS error changes with changing time of evolution for different numbers of training states using ITE and ASP respectively.
@@ III.1 Application to 5-site XY spin chain
We demonstrate both these approaches for a test model: a 5-site XY open spin chain with mixed magnetic fields,
$\mathcal{H}=$ $J\sum_{i=1}^{N-1}\left(X_{i}X_{i+1}+Y_{i}Y_{i+1}\right)$
$+B_{Z}\sum_{i=1}^{N}Z_{i}+B_{X}\sum_{i=1}^{N}X_{i}$ (5)
where we fix $N=5$, $J=1$ and $B_{X}=0.2$, and vary $B_{Z}\in[0,3]$. A finite transverse field in the $X$ direction ($B_{X}$) opens up an energy gap at level crossings — this is necessary to lift the degeneracy at level crossings, allowing adiabatic time evolution to follow the lowest energy state (see Fig. A3). In the next section, we will address these issues in detail. For our analysis, we will consider $20$ equally distributed points in varying $B_{Z}$ range as target Hamiltonians. For those points, we calculate the ground state energy eigenvalues solving the GEP (Section II) with the subspace formed by truncated ITE/ASP states. To compare errors with different levels of truncation for both methods, we will consider RMS error for all the target points, given by:
@@
In Fig. 2 we demonstrate how using eigenvector continuation improves both ITE and ASP ground state preparation algorithms. We apply truncated versions of both algorithms to the 5-site XY model, using a modest number of time steps (40 and 75 for ITE and ASP, respectively) to get the subspace basis vectors. For comparison, obtaining a reasonable ground state would require 600 and 750 steps for ITE and ASP respectively, for all of the target points. To get the basis vectors, we choose equally spaced $k_{p}=5$ training points for the energy spectrum and fidelity analysis and $k_{p}=\{3,4,5\}$ training points for the point of truncation analysis.
@@
Moving to truncated ASP, we again show the spectrum together with the obtained energies from truncated ASP in panel (d). Here, we have used $T_{\mathrm{max}}=3.75$, and $dt=0.05$, i.e. 75 time steps, and we use a linear ramp in $B_{z}$ with $dB_{z}/dt=0.8$. The process starts at $B_{Z}=3$, where the ground state is close to the fully polarized state, and thus is easy to prepare. The protected level crossings are turned into avoided crossings by the $B_{X}=0.2$ which induces gaps far smaller than would be appropriate for this particular ramp. Not surprisingly, the truncated ASP does not manage to follow the ground state due to early truncation of the evolution, as the Hamiltonian is changing too rapidly for the adiabatic theorem to hold. Nevertheless, we can use the truncated ATE states to produce a subspace, in which the ground state is accurately captured. The resulting EC spectrum and fidelity are shown in panels (d) and (e), respectively, where we can observe that EC produces a marked improvement in both quantities.
@@
To make a quantitative comparison, we consider the RMS error ($\Delta\varepsilon_{rms}$) between the exact spectrum and the EC spectrum for the 20 equally spaced target points under consideration (Eq. 6) using both methods. We consider the RMS error as a function of $\tau_{\mathrm{max}}$ for ITE, which accounts for the total evolution time, and $T_{\mathrm{max}}$ for ATE, which accounts for the rate of change of the Hamiltonian. In both cases, the use of EC significantly improves the approximate results. At 5 training points (indicated by dashed vertical lines), which corresponds to the spectra and fidelities shown in Fig. 2), we find a 78% and 97% reduction in error for ITE and ASP, respectively. The RMS error generally continues to reduce as a function of $\tau_{\mathrm{max}}$/$T_{\mathrm{max}}$. One notable exception occurs when $k_{p}=3$, where the error increases for some values of $\tau_{\mathrm{max}}$/$T_{\mathrm{max}}$. This illustrates the use of choosing training points that do not lie in parameter regions that are difficult to calculate — with 3 training points, one of the critical parameters is right at an avoided level crossing, which leads to the increase in the RMS error. This issue can readily be solved by adding more training states.
@@
For a comparison of resources required, let us first consider ASP. To obtain an accurate ground energy spectrum for a set of target Hamiltonians, using ASP we require very slow (adiabatic) evolution. In our example, we would require $750$ time steps to get the fidelity close to $1$ (see Fig. A2). Using truncated ASP together with EC, using only $75$ time steps to get the basis vectors is sufficient (c.f. Fig. 2 (f)). Similarly, using ITE, the total time of evolution required for the desired error is $\tau_{\mathrm{max}}=6$, i.e. $30$ imaginary time steps at each target point, yielding $30\times 20=600$ total steps. With truncation, we used $\tau_{\mathrm{max}}=1.6$ which is $8$ time steps for $k_{p}=5$ points, requiring $40$ total steps to get the basis vectors. After getting the projected Hamiltonian and overlap matrix, all the energy values are calculated by solving GEP, which is classically easy given matrix sizes in GEP depend on the number of basis vectors. Overall, using EC on top of these standard state preparation techniques saves $>90\%$ of the time evolution steps.

### v3

@@ Abstract
Subspace methods are powerful, noise-resilient methods that can effectively prepare ground states on quantum computers. The challenge is to get a subspace with a small condition number that spans the states of interest using minimal quantum resources. In this work, we will use eigenvector continuation (EC) to build a subspace from the low-lying states of a set of Hamiltonians. The basis vectors are prepared using truncated versions of standard state preparation methods such as imaginary time evolution (ITE), adiabatic state preparation (ASP), and variational quantum eigensolver (VQE). By using these truncated methods combined with eigenvector continuation, we can directly improve upon them, obtaining more accurate ground state energies at a reduced cost. We use several spin systems to demonstrate convergence even when methods like ITE and ASP fail, such as ASP in the presence of level crossings and ITE with vanishing energy gaps. We also showcase the noise resilience of this approach beyond the gains already made by having shallower quantum circuits. Our findings suggest that eigenvector continuation can be used to improve existing state preparation methods in the near term.
@@ I Introduction
Finding the ground state of non-integrable systems is one of the primary targets of studies in quantum mechanics. One can achieve these goals via classical numerical methods, such as exact diagonalization or tensor network algorithms. However, these methods have limitations, mainly for two- and three-dimensional systems and/or for highly entangled states [1, 2]. As a result, there has been significant interest in quantum computing alternatives to address this issue. Within the setting of quantum computing, the problem is recast as finding an efficient way to prepare a desired quantum state. This problem is generally QMA-hard [3, 4, 5], i.e. difficult even for quantum computers; nevertheless, a significant research effort is being directed towards making inroads into this important problem, as quantum devices are believed to be able to approach this problem and perform much better than their classical counterparts [6, 7]. This promise is not yet realized, and problems such as barren optimization plateaus [8, 9, 10] plague the current methods. While fault tolerant algorithms that have provable convergence rates do exist, these are not yet realizable on near-term hardware, which is noisy and limited in its capabilities. Because of this reason, researchers are focusing on developing NISQ-friendly algorithms that can still perform well with hardware imperfections [11, 12, 13].
Several quantum state preparation algorithms can be categorized into these broad categories, including variational quantum algorithms (e.g, VQE, QAOA) [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], time evolution methods (e.g, ASP, QITE) [24, 25, 26, 27, 28, 29], and subspace methods (e.g, Krylov subspace) [30, 31, 32, 33, 34, 35]. While they all have advantages and disadvantages, none are ideal for implementation on current hardware. Variational quantum algorithms use classical optimization-based and learning-based approaches with parameterized quantum circuits but face optimization landscape issues along with trainability, accuracy, and efficiency limitations [8]. Time evolution methods, and in particular adiabatic state preparation (ASP), provide an efficient mapping to a quantum circuit; however, ASP often requires a long evolution [36, 37, 26], an easily prepared starting state, and cannot be used in the presence of symmetry-protected level crossings. Imaginary time evolution [38] (ITE) guarantees convergence, but implementing non-unitary evolution in a quantum circuit is complicated. Both evolution-based algorithms tend to involve deep quantum circuit implementations; and as the system size and the corresponding Hilbert space grow, these algorithms require a concomitantly longer circuit to converge [39, 40, 41, 42], thus increasing the quantum resources required.
@@
In this work, we will improve on this idea by realizing that an exact ground state is not needed for EC to work; rather, obtaining a number of states with a decent overlap with the low energy manifold suffices. Here we will use real- and imaginary-time ground-state preparation techniques, although other methods could be used similarly to get the low-energy states. In all cases, an otherwise insufficiently deep (or truncated) circuit can be used. Getting low-energy states is much easier and more cost-effective than obtaining exact ground states; we will explore these ideas for time evolution methods in the next section. We will show how this method can be used to handle issues with time-evolution methods like vanishing energy gaps and/or level crossings, and saving resources in doing so. A similar idea of using the approximate state preparation for getting a basis for subspace diagonalization was explored in a recent work [53], where they use approximate VQE for preparing a subspace to solve quantum chemistry instances, but remain focused on a single Hamiltonian.
The outline of the paper is as follows: In section II, we will outline the method of eigenvector continuation. In section III, we will apply it using time evolution and variational state preparation methods to get a low-energy subspace for the XY spin model. In section IV, we will extend the application of the method to address the issue of vanishing energy gaps using the time-evolution methods with an example of a model with level crossings, XY model, and model with changing degeneracies in the ground state, Kagome XXZ model. Section V discusses the noise resilience of the method by performing noisy simulations. Lastly, we conclude with a discussion in section VI.
@@ II Eigenvector Continuation
In this work, we still use a set of basis states that spans the low-energy subspace, but we propose using state preparation methods that are truncated before they reach the ground state. This gives us a low-energy state that is not necessarily an eigenstate or the ground state, rather, this is a superposition of a few low-energy states of the system. As we will show, getting these states from the small number of training points gets us enough overlap with all the ground states to be able to obtain reasonable ground states from the generalized eigenvalue problem.
@@
For the training Hamiltonians, we find approximate ground states, using any state preparation method, to obtain a set of low-energy vectors that will form the basis vectors for a subspace $(\phi_{t}(\theta)=\{\ket{\phi_{1}},\ket{\phi_{2}},\dots\ket{\phi_{k_{p}}}\})$. In this paper we have focused on time evolution methods for state preparation—ASP and ITE; but we also provide one example using a variational method to showcase the versatility of our approach.
Now that we have a subspace that spans low energy states, we project the target Hamiltonians $(H(\theta_{i}))$ onto this subspace and solve the generalized eigenvalue problem:
@@
One can solve Eq. (II) via standard eigendecomposition routines. Ill-conditioned cases occur whenever $\mathbb{S}$ is non-positive definite, thus requiring a separate diagonalization of $\mathbb{S}$, followed by applying some truncation scheme for the corresponding eigenpairs and a basis rotation of $\mathbb{H}$ [43, 46]. These extra steps were unnecessary in this work, as we have a positive-definite overlap matrix for all the cases studied here.
The lowest eigenvalue thus obtained gives the ground state energy of the target Hamiltonian. The overlap matrix $(\mathbb{S})$ of the basis vectors remains the same for all the target points, but the projected Hamiltonian ($\mathbb{H}$) needs to be calculated separately for each target point. We can use the fact that the target Hamiltonians differ only by the values of the coefficients of the Hamiltonian operators and simplify the process of measuring the projected Hamiltonians ($\mathbb{H}$), measuring all the Hamiltonian operators separately and adding them with corresponding coefficients for each target point.
@@ III Improving Quantum state preparation methods using EC
In this work, we will use time evolution-based methods, i.e. adiabatic state preparation (ASP) and imaginary time evolution (ITE), for generating our low-energy subspace. For completeness, we also show one example using the variational quantum eigensolver (VQE) to obtain the training points. However, contrary to time evolution methods, there is no straightforward way to truncate VQE; therefore, we dedicate most of our presentation to results obtained via ASP and ITE.
To improve the quality of the state prepared via time evolution methods, one can consider either a longer evolution time, or a smaller time step; we will focus on the former. To illustrate the use of EC to improve a truncated state preparation, we truncate at $\leq 10\%$ of the time evolution steps required to get reasonable ground state energies.
In what follows, we briefly describe the state preparation methods considered here.
@@ III.0.2 Adiabatic State Preparation
III.0.3 Variational Quantum Eigensolver
The Variational Quantum Eigensolver (VQE) is one of the most promising methods for ground-state preparation in the NISQ era. It stems from the variational principal of quantum mechanics, which states that the expectation value of a Hamiltonian $\mathcal{H}$ with respect to a generic parameterized state $\left|\psi\left(\bm{\theta}\right)\right\rangle$ is bounded from below by the Hamiltonian ground-state energy $E_{0}$, i.e.,
$E_{0}\leq\left\langle\psi\left(\bm{\theta}\right)\right|\mathcal{H}\left|\psi\left(\bm{\theta}\right)\right\rangle$ (4)
Therefore, by defining a parameterized ansatz that builds $\left|\psi\left(\bm{\theta}\right)\right\rangle$, one can construct the corresponding quantum circuit and measure the energy, which is fed into a classical minimizer routine as the cost function. The minimizer then returns a new set of optimized parameters $\bm{\theta}$. One repeats this procedure until the classical minimizer is converged, thus yielding an approximated ground state of the underlying system described by $\mathcal{H}$.
In general VQE applications, one builds the state $\left|\psi\left(\bm{\theta}\right)\right\rangle$ by applying many layers of a predefined unitary operator $U\left(\bm{\theta}\right)$ to a classical product state $\left|\psi_{0}\right\rangle$, i.e.,
$\left|\psi\left(\bm{\theta}\right)\right\rangle=\prod_{\ell}U\left(\bm{\theta}_{\ell}\right)\left|\psi_{0}\right\rangle$ (5)
However, selecting the optimal ansatz $U\left(\bm{\theta}\right)$ is a problem-dependent question that has led to many different proposals of VQE ansatze [17]. In this work we choose the one known as Hamiltonian Variational Ansatz (HVA), which, as the name suggests, takes the operators that appear in the system Hamiltonian to build the variational state [58, 59]. We explicitly show the ansatz considered here in the next section, after the system Hamiltonian is introduced.
In addition, differently from ASP and ITE, VQE does not have a single variable to be chosen as truncation parameter; for instance, one could choose to truncate the number of variational parameters $\bm{\theta}$, or the number of ansatz layers, or the number of iterations within the classical optimizer. In this work we choose the latter as our truncation parameter. We also use the maximum number of variational parameters (i.e., one per gate) and keep the number of HVA layers as low as possible.
@@ III.1 Application to 5-site XY spin chain
We demonstrate these approaches for a test model: a 5-site XY open spin chain with mixed magnetic fields,
$\mathcal{H}=$ $J\sum_{i=1}^{N-1}\bigg{[}X_{i}X_{i+1}+Y_{i}Y_{i+1}\bigg{]}$
$+\sum_{i=1}^{N}\bigg{[}B_{Z}Z_{i}+(-1)^{i}B_{X}X_{i}\bigg{]}$ (6)
where we fix $N=5$, $J=1$ and a staggered $B_{X}=0.2$, and vary $B_{Z}\in[0,3]$. A finite staggered transverse field in the $X$ direction ($B_{X}$) opens up an energy gap at level crossings — this is necessary to lift the degeneracy at level crossings, allowing adiabatic time evolution to follow the lowest energy state (see Fig. A3). In the next section, we will address these issues in detail.
Hence, given the Hamiltonian in Eq. (III.1), one layer of the HVA ansatz corresponds to
$U\left(\bm{\theta}\right)=U_{YY}\left(\bm{\delta}\right)U_{XX}\left(\bm{\gamma}\right)U_{Z}\left(\bm{\beta}\right)U_{X}\left(\bm{\alpha}\right)$ (7)
where
$U_{YY}\left(\bm{\delta}\right)$ $=\textrm{exp}\left(-i\sum_{\left\langle i,j\right\rangle}\delta_{i,j}Y_{i}Y_{j}\right)$
$U_{XX}\left(\bm{\gamma}\right)$ $=\textrm{exp}\left(-i\sum_{\left\langle i,j\right\rangle}\gamma_{i,j}X_{i}X_{j}\right)$
$U_{Z}\left(\bm{\beta}\right)$ $=\textrm{exp}\left(-i\sum_{i}\beta_{i}Z_{i}\right)$
$U_{X}\left(\bm{\alpha}\right)$ $=\textrm{exp}\left(-i\sum_{i}\alpha_{i}X_{i}\right)$
Fig. 2 depicts the quantum circuit representation of Eq. (7). For the simulations of the 5-site XY model, we have considered an ansatz with two HVA layers, with each layer following the same operator ordering as shown in Fig. 2.
Figure 2: Schematic of a single layer of the Hamiltonian Variational Ansatz (HVA) for a 5-site XY model with transverse and longitudinal external fields.
For our analysis, we will consider $20$ equally distributed points in varying $B_{Z}$ range as target Hamiltonians. For those points, we calculate the ground state energy eigenvalues solving the GEP (Section II) with the subspace formed by truncated states. To compare errors with different levels of truncation, we will consider RMS error for all the target points, given by:
@@
Figure 3: Implementing EC with truncated ASP, ITE, and VQE for a 1D chain XY model 5 site. Panels (a), (d), and (g) show the system complete energy spectrum (full lines) as a function of the longitudinal field $B_{z}$, with the ground-state energy curve highlighted in black. The green diamonds represent the reference points we used for building the subspace and computing the corresponding EC curve (red stars). The arrows indicate the direction taken by each state preparation method. Panels (b), (e), and (h) compare the fidelities of EC and truncated state preparation data with respect to the exact ground state. Panels (d), (f), and (i) show the energy error [Eq. (8)] as a function of truncation parameters, which are evolution time for ITE and ASP and number of iterations in the minimizer routine for VQE. Here we have considered three different subspace dimensions $k_{p}$ for obtaining the EC data, which we also compare to the truncated state preparation results without EC (blue dashed curve). The vertical lines represent the truncation parameter value used for the data shown in the other panels.
In Fig. 3 we demonstrate how using eigenvector continuation improves the ground state preparation algorithms. For the time evolution methods, we considered a modest number of time steps (40 and 75 for ITE and ASP, respectively) to get the subspace basis vectors. For comparison, obtaining a reasonable ground state would require 600 and 750 steps for ITE and ASP respectively, for all of the target points. For the truncated VQE, we considered 12 iterations of the classical optimizer for all the training points. For comparison, the unbounded minimizer would require from 73 to 869 iterations to converge for these exact points. To get the basis vectors, we choose equally spaced $k_{p}=5$ training points for the energy spectrum and fidelity analysis and $k_{p}=\{3,4,5\}$ training points for the point of truncation analysis.
@@
Moving to truncated ASP, we again show the spectrum together with the obtained energies from truncated ASP in panel (d). Here, we have used $T_{\mathrm{max}}=3.75$, and $dt=0.05$, i.e. 75-time steps, and we use a linear ramp in $B_{z}$ with $dB_{z}/dt=0.8$. The process starts at $B_{Z}=3$, where the ground state is close to the fully polarized state, and thus is easy to prepare. The protected level crossings are turned into avoided crossings by the $B_{X}=0.2$ which induces gaps far smaller than would be appropriate for this particular ramp. Not surprisingly, the truncated ASP does not manage to follow the ground state due to early truncation of the evolution, as the Hamiltonian is changing too rapidly for the adiabatic theorem to hold. Nevertheless, we can use the truncated ASP states to produce a subspace, in which the ground state is accurately captured. The resulting EC spectrum and fidelity are shown in panels (d) and (e), respectively, where we can observe that EC produces a marked improvement in both quantities.
Finally, panels (g) and (h) show the energies and fidelities calculated from the truncated VQE subspace. In this case, the selected training points (i.e., green diamonds) are closer to the exact ground state curve than the other two methods. Nevertheless, similarly to ITE and ASP, we observe a considerable improvement in the fidelities after we perform the subspace expansion, with the minimum fidelity for the considered $B_{z}$ range jumping from 21% (blue dashed curve) to 98% (red stars); a significant gain with little computational cost.
@@
To make a quantitative comparison, we consider the RMS error ($\Delta\varepsilon_{rms}$) between the exact spectrum and the EC spectrum for the 20 equally spaced target points under consideration (Eq. 8) using both methods. We consider the RMS error as a function of $\tau_{\mathrm{max}}$ for ITE, which accounts for the total evolution time, $T_{\mathrm{max}}$ for ASP, which accounts for the rate of change of the Hamiltonian, and the number of iterations of the classical optimizer for VQE. For all these methods, one can notice that the use of EC significantly improves the approximate results. At 5 training points, which corresponds to the spectra and fidelities shown in Fig. 3, we find a 78% and 97% reduction in error for ITE and ASP, respectively. The RMS error generally continues to reduce as a function of $\tau_{\mathrm{max}}$/$T_{\mathrm{max}}$. One notable exception occurs when $k_{p}=3$, where the error increases for some values of $\tau_{\mathrm{max}}$/$T_{\mathrm{max}}$. This illustrates the use of choosing training points that do not lie in parameter regions that are difficult to calculate — with 3 training points, one of the critical parameters is right at an avoided level crossing, which leads to the increase in the RMS error. This issue can readily be solved by adding more training states.
@@
For a comparison of resources required, let us first consider ASP. To obtain an accurate ground energy spectrum for a set of target Hamiltonians, using ASP we require very slow (adiabatic) evolution. In our example, we would require $750$ time steps to get the fidelity close to $1$ (see Fig. A2). Using truncated ASP together with EC, using only $75$ time steps to get the basis vectors is sufficient (c.f. Fig. 3 (f)). Similarly, using ITE, the total time of evolution required for the desired error is $\tau_{\mathrm{max}}=6$, i.e. $30$ imaginary time steps at each target point, yielding $30\times 20=600$ total steps. With truncation, we used $\tau_{\mathrm{max}}=1.6$ which is $8$ time steps for $k_{p}=5$ points, requiring $40$ total steps to get the basis vectors. After getting the projected Hamiltonian and overlap matrix, all the energy values are calculated by solving GEP, which is classically easy given matrix sizes in GEP depending on the number of basis vectors. Overall, using EC on top of these standard state preparation techniques saves $>90\%$ of the time evolution steps.